OpenStudy (anonymous):

Medal for the first person to solve this! Choose the equation below that represents the line that passes through the point (7, −2) and has a slope of −3. y − 2 = −3(x + 7) y + 7 = −3(x − 2) y + 2 = −3(x − 7) y − 7 = −3(x + 2)

Vocaloid (vocaloid):

use point slope formula y - y1 = m(x-x1) where m = slope, x1 = 7, and y1 = -2

OpenStudy (anonymous):

Frieza is my favorite super hero of all time

OpenStudy (anonymous):

Vocaloid your answer has me confused can you further explain

OpenStudy (alexandervonhumboldt2):

@ryandaman2017 please do not post spam comments.

Vocaloid (vocaloid):



OpenStudy (anonymous):

lIt was spam tho i didnt mind his comment either

OpenStudy (anonymous):

thanks Vocaloid take this medal

OpenStudy (anonymous):

how was that spam? I was just saying because I was interested in his username

OpenStudy (alexandervonhumboldt2):

posting comments irrelevant to the post is spamming. you can say such things in private messages.